# Limits to Predicting Online Speech Using Large Language Models

### v1

@@ Abstract
We study the predictability of online speech on social media, and whether predictability improves with information outside a user’s own posts. Recent work suggests that the predictive information contained in posts written by a user’s peers can surpass that of the user’s own posts. Motivated by the success of large language models, we empirically test this hypothesis. We define unpredictability as a measure of the model’s uncertainty, i.e., its negative log-likelihood on future tokens given context. As the basis of our study, we collect a corpus of 6.25M posts from more than five thousand X (previously Twitter) users and their peers. Across three large language models ranging in size from 1 billion to 70 billion parameters, we find that predicting a user’s posts from their peers’ posts performs poorly. Moreover, the value of the user’s own posts for prediction is consistently higher than that of their peers’. Across the board, we find that the predictability of social media posts remains low, comparable to predicting financial news without context. We extend our investigation with a detailed analysis about the causes of unpredictability and the robustness of our findings. Specifically, we observe that a significant amount of predictive uncertainty comes from hashtags and @-mentions. Moreover, our results replicate if instead of prompting the model with additional context, we finetune on additional context.
1 Introduction
We revisit the problem of predicting online speech in light of dramatic advances language modeling. We focus on the central question: How predictable is our online speech using large language models? Such predictive capabilities could inform research on substantial risks – such as user profiling, impersonation and exerting influence on real users Carroll et al. (2023); Weidinger et al. (2021). Inspired by work exploring the possibility of user profiling Bagrow et al. (2019), we explore the question of predicting user speech from their peers. We ask the following: How predictable is a social media post given posts from the author’s peers? We contrast the answer with how predictable social media posts are from the authors’ own posts. Through an experiment spanning millions of tweets and thousands of subjects, our study provides a detailed picture of the current state of predicting online speech.
@@ 1.1 Contributions
We investigate the predictability of online speech on the social media platform X (Twitter) using a corpus of 6.25M posts (tweets) of 5000 subjects and their peers. We use three large language models of increasing size, GPT-2-XL-1.5B, Falcon-40B, and Llama-2-70B. We use these models to estimate the predictability of our subjects’ posts under various settings.
Predictability improves significantly with model size, as expected. However, even the best model performs poorly without additional context. For illustration, most users’ posts are less predictable than financial news. We therefore test how much prediction benefits from three additional sources of context, specifically:
@@
To understand the sources of unpredictability, we further investigate the locations of greatest uncertainty within a post. We find that @-mentions and hashtags are typically the points of greatest uncertainty within a post. Consequently, removing both generally improves predictability across the board. The relative comparisons, however, stay unchanged.
Large language models can be sensitive to prompting strategies. To test the robustness of our findings, we replace prompting with fine-tuning. Specifically, instead of prompting GPT-2-XL with a specific source of context, we fine-tune the model on the same source of context. We find that the relative comparisons between user, peer, and random context, remain the same.
@@ 2 Related work
Understanding and modeling language usage on social media Bashlovkina et al. (2023); Kern et al. (2016); Schwartz et al. (2013) has become of particular interest to the research community, where Twitter is one of the most studied social media platforms Zhang et al. (2023); Qudar and Mago (2020). The most common method is using large language models (LLMs), where they finetune on social-media data and evaluate models on specific downstream tasks such as sentiment analysis, hate speech detection, etc. Zhang et al. (2023); Bashlovkina et al. (2023); Qudar and Mago (2020) We are interested in a more general notion of predictability – in particular, the predictability of individual users’ online speech. Following common practice in natural language processing, we take model uncertainty as a measure of how hard it is for the model to predict some text.
@@ 3 Experimental setup
We consider a similar experimental setup as Bagrow et al. (2019), but with recent large language models and 15 peers per subject. We define predictability (or rather unpredictability) as a measure of the model’s uncertainty, i.e., its negative log-likelihood, on a specific user’s tweets conditional on given context. We collected tweets from a sample of $\sim$5k subjects. We experiment with three types of contexts: (i) past tweets from the user (or “user context” - Fig. 2(b)), (ii) past tweets from the user’s social circle (“peer context”) and finally (iii) past tweets from a control group (“control” or “random context”).
We start by describing our data collection process (Section 3.1) and what models we used (Section 3.2). Then, we go into detail about the implementation of our prompting and finetuning experiments (in Section 3.3 and 3.4, respectively).
@@
Figure 3: Our data collection process can be divided into two stages. In the first stage (Fig. 3(a)), we collected 10M tweets in early 2023 which served as our base for sampling subjects. In the second stage (Fig. 3(b)), we collected users’ timelines. First, we collected timelines of subjects. Based on those, we determined which users they had most frequently mentioned (their “peers”). We selected the top-15 most mentioned peers and 15 random users per subject, and collected their timelines too.
@@ 3.1 Data Collection
The data collection process consisted of two main stages: an initial sampling period where we recorded real-time Twitter activity for a month in early 2023 and a second stage where we collected the timelines of sampled users. For a high-level overview of the data collection process, see Figure 3. We collected tweets from three groups of users:
@@
    subjects, approximately 5,000 randomly sampled Twitter users,
@@
    peers of subjects, which we take to be the top 15 people that each subject most frequently mentions,
  3. 3.
    random users for control purposes.
Throughout our data collection, we only collected tweets that were written by the user who posted them (e.g. no retweets) and were classified as English according to Twitter’s own classification algorithm.
3.1.1 Sampling stage
Sampling was done by collecting a pool of tweets whose authors would serve as our base for picking our subjects. We collected them using the Twitter Firehose API; which allowed us to collect a 1% sub-sample of real-time tweet activity. This collection period lasted roughly 30 days and was done in early 2023 (from 20. January to 10. February), during which we collected 10M tweets (with $\sim$5M unique authors). We sampled $N=5102$ subjects from this pool ($0.1\%$ of authors) for our experiments. Twitter had roughly $\sim$500M active monthly users in 2023 Yaccarino (2023) – some of which may have been bot accounts Lee et al. (2011). To address this issue, we decided to filter out accounts that scored high (above 0.5 on a scale of 0 to 1) on the Bot-O-Meter bot detection API Sayyadiharikandeh et al. (2020). Additionally, we also filtered out users that had a high retweet ratio, where $>80\%$ of their feed consisted of retweets. This ensured the selection of users who had a sufficient amount of self-authored tweets for our dataset.
3.1.2 Timeline collection stage
For each subject, we collected a total of 500 tweets $\mathcal{T}_{u}=\mathcal{T}^{\texttt{eval}}_{u}\cup\mathcal{T}^{\texttt{user}}_{u}$ from their timelines. Half of those tweets were used for estimating predictability $\mathcal{T}^{\texttt{eval}}_{u}$, while the other half $\mathcal{T}^{\texttt{user}}_{u}$ served as context. Some tweets in $\mathcal{T}_{u}$ date back as early as 2011, however the vast majority (95%) were authored after 2022. Collected tweets were authored no later than the start of our sampling stage (20. January 2023).
Besides user context, we also introduce “peer” and “random” context: $\mathcal{T}^{\texttt{peer}}_{u}$ and $\mathcal{T}^{\texttt{random}}_{u}$, again with 250 tweets each. We selected the 250 most recent tweets from a pool of 15 users’ timelines. Peers were the top-15 users who were most frequently mentioned by a subject. Random users were sampled from a total of $\sim$90k timelines that we had queried in this stage. We always made sure that context tweets were authored before the oldest tweet inside $\mathcal{T}^{\texttt{eval}}_{u}$. In total, we collected approximately 6.25M tweets for our experients.
Similarly to Bagrow et al. (2019), we created a second control group (temporal control). However, because of the similarity of the results, we only report results on the random control group (social control) in the main text. For results including the temporal control, please see Appendix B.4. For more detailed information on the dataset and how it was collected, we refer the reader to Appendix A.1. There we go into detail on what filters we used to achieve a high-quality, representative dataset of English tweets and share some statistics about our subject pool.
3.2 Models
We selected three large language models for our experiments: GPT-2-XL with 1.5B parameters Radford et al. (2019), Falcon with 40B parameters Almazrouei et al. (2023) and Llama-2 with 70B parameters Touvron et al. (2023). They represent models that share a similar transformer-based architecture that have recently become popular due to their generative capabilities. However, they differ in parameter size (1.5B, 40B and 70B parameters, respectively), training corpus and tokenizers. We considered using GPT-3 and GPT-4, however the OpenAI API unfortunately does not allow access to log probabilities for all tokens (which is necessary for our analysis), only to the top-5. We used Huggingface’s transformers library Wolf et al. (2020) to load the models and run our experiments. For the prompting experiments, we loaded the larger models using 8-bit precision (which has little to no impact on performance Dettmers et al. (2022)) so that they could fit on 2 A-100 80GB GPUs. We only use GPT-2-XL for our finetuning experiments.
@@ 3.3 Prompting Experiments
Let us denote a single tweet as $T=(t_{1},t_{2},...t_{m})$, where $t_{i}$ is a single token ($i=0...m$). A token is an item in the LLM’s vocabulary, and can be thought of as a collection of characters that frequently co-occur. Each tweet has a maximum of 280 characters [^1] This limit changed to 4000 characters on the 09. February 2023. and after tokenization most tweets have between 0-100 tokens (Fig. 16 in Appendix). We use a language model with parameters $\theta$ to predict token $t_{i}$ based on the preceding tokens $t_{<i}$. The model’s output will be a likelihood over all possible tokens in the model’s vocabulary, however we are only interested in the conditional probability of $t_{i}$: $p_{\theta}(t_{i}|t_{<i})$.
3.3.1 Calculating negative log-likelihood
@@
where $n$ is the total number of tokens. We additionally introduce notation to distinguish what context was used to calculate: $\bar{L}_{u}^{c}$, where subscript $c\in\{\texttt{user},\texttt{peer},\texttt{random}\}$. Here, the conditional probability of token $t_{i}$ is based on preceding tokens $t_{<i}$ as well as tokens from the appropriate context: $p_{\theta}(t_{i}|\mathcal{T}_{u}^{c},t_{<i})$. The added context lends a convenient cross-entropy like interpretation of the shared information between the context and evaluation tweets we are trying to predict. Tweets that served as context were concatenated using the ‘newline’ token for GPT-2-XL and the ‘space’ token for the two bigger models (they had no standalone ‘newline’ token like GPT-2). We used the maximum input sequence length available for each model (1024 tokens for GPT-2-XL, 2048 for Falcon 40B and 4096 for Llama-2) to calculate the token probabilities. In case the provided context exceeded this length, the oldest tweets were discarded.
3.3.2 Conversion to bits per character
@@
3.3.3 Estimating the effect size of context on predictability
@@ 3.4 Finetuning Experiments
In this section we present a second method for estimating predictability, which has two advantages over our previous prompting approach. First, we can fit the entire context into the model compared to prompting where the maximum size of the input sequence was a limitation. Second, it bypasses a common ailment of LLMs: their sensitivity to prompting strategy. However, finetuning can also be very expensive—which is why we decided to limit our experiments to GPT-2-XL only.
Tweets were concatenated (with the special eos token ’<|endoftext|>’ serving as a separator between them) then divided into chunks of 1024 tokens so that they fit into the model. Here, our proxy for unpredictability was the negative log-likelihood (or also commonly known as the cross-entropy loss) in the final round of fine-tuning. Reported results are calculated on $\mathcal{T}^{\texttt{eval}}_{u}$.
Our finetuning experiments had two stages. In the first pre-finetuning stage, we trained on a large set of tweets to condition the model to learn basic Twitter syntax and common vocabulary. This later helped us learn user related features much more efficiently in the second stage. After pre-finetuning, we additionally finetuned on one of the contexts. We also experimented with using mixtures of different contexts. We used the example finetuning script from the transformers library as our base333run_clm.py from here: https://github.com/huggingface/transformers/tree/main/examples/pytorch/language-modeling, where we kept most of the default training arguments. Some finetuning choices were inspired by an open-source project called HuggingTweets Dayama (2022).
3.4.1 Pre-finetuning on 10M tweets
We finetuned GPT-2-XL on the 10M tweets we collected during the sampling stage of our data collection process (as described Section 3.1), where a 5% split was reserved for validation. We finetuned all parameters for 1 epoch, using a constant learning rate of $5\mathrm{e}{-5}$, batch size 8 and fp16 mixed precision training on a single A-100 80GB GPU. More details on the pre-finetuning and loss curves can be found in Appendix C.1.
3.4.2 Finetuning on context
Starting from the pre-finetuned model, for each user we finetuned a model on $\mathcal{T}_{u}^{c}$, for each $c\in\{\texttt{user},\texttt{peer},\texttt{random}\}$. That is three separate models per subject, resulting in a total of $\sim$15k finetuned models. We finetuned for 5 epochs with constant learning rate $1\mathrm{e}{-5}$ and batch size 1. We tracked the cross-entropy loss on $\mathcal{T}^{\texttt{eval}}_{u}$ periodically.
We also include experiments on finetuning on a mixture of contexts (eg. peer+random containing tweets from both $\mathcal{T}^{\texttt{peer}}_{u}$ and $\mathcal{T}^{\texttt{random}}_{u}$). We combined them by sampling an equal amount of tweets uniformly from each context, while keeping the total number of tweets the same (250 tweets) to enable a fair comparison. Again, the base model was our pre-finetuned model, and we finetuned for 5 epochs with constant learning rate $1\mathrm{e}{-5}$ and batch size 1.

### v2

@@ Abstract
We study the predictability of online speech on social media, and whether predictability improves with information outside a user’s own posts. Recent theoretical results suggest that posts from a user’s social circle are as predictive of the user’s future posts as that of the user’s past posts. Motivated by the success of large language models, we empirically test this hypothesis. We define predictability as a measure of the model’s uncertainty, i.e., its negative log-likelihood on future tokens given context. As the basis of our study, we collect 10M tweets for “tweet-tuning” base models and a further 6.25M posts from more than five thousand X (previously Twitter) users and their peers. Across four large language models ranging in size from 1.5 billion to 70 billion parameters, we find that predicting a user’s posts from their peers’ posts performs poorly. Moreover, the value of the user’s own posts for prediction is consistently higher than that of their peers’. We extend our investigation with a detailed analysis on what’s learned in-context and the robustness of our findings. From context, base models learn to correctly predict @-mentions and hashtags. Moreover, our results replicate if instead of prompting the model with additional context, we finetune on it. Across the board, we find that predicting the posts of individual users remains hard.
@@
1 Introduction
We revisit the problem of predicting online speech in light of dramatic advances in language modeling. We focus on the central question: How predictable is our online speech using large language models? Such predictive capabilities could inform research on substantial risks – such as user profiling, impersonation and exerting influence on real users Carroll et al. (2023); Weidinger et al. (2021). Inspired by work exploring the possibility of user profiling through their peers Bagrow et al. (2019), we ask the following: How predictable is a social media post given posts from the author’s peers? We contrast the answer with how predictable social media posts are from the authors’ own posts. Through experiments spanning millions of tweets and thousands of subjects, our study provides a detailed picture of the current state of predicting online speech.
@@ 1.1 Contributions
We investigate the predictability of online speech on the social media platform X (Twitter) using a corpus of 6.25M posts (tweets) of 5000 subjects and their peers. An additional 10M tweets is reserved for tweet-tuning models. We use four large language models of increasing size, GPT-2-XL-1.5B, Llama-3-8B, Falcon-40B, and Llama-2-70B. We use these models to estimate the predictability of our subjects’ posts under various settings.
Predictability improves significantly with model size, as expected. However, even the best model performs poorly without additional context. For illustration, most users’ posts are less predictable than posts from financial news accounts (Fig. 9). We therefore test how much prediction benefits from three additional sources of context, specifically:
@@
To understand what is being learned in-context, we further investigate the locations of greatest uncertainty within a post. We find that base models learn to correctly predict @-mentions and hashtags. Consequently, removing both generally improves predictability across the board. The relative comparisons, however, stay unchanged.
Large language models can be sensitive to prompting strategies. To test the robustness of our findings, we replace prompting with finetuning. Specifically, instead of prompting a tweet-tuned model with a specific source of context, we finetune the model on the same source of context. We find that the relative comparisons between user, peer, and random context, remain the same.
@@ 2 Related work
Understanding and modeling language usage on social media Bashlovkina et al. (2023); Kern et al. (2016); Schwartz et al. (2013) has become of particular interest to the research community, where Twitter is one of the most studied social media platforms Zhang et al. (2023); Qudar and Mago (2020). Feeding user generated content to language models could serve both as a mirror and a tool to understand and (potentially) steer users more effectively. For example, LLMs can be used to infer / extract sensitive author information from what they post online Staab et al. (2023). They can also be leveraged to predict users’ political leaning, morality or toxicity Jiang and Ferrara (2023). Without focusing on any particular downstream task, we are interested in how well these models are able to model individual users’ online speech. In other words, how well models are able to learn the distribution of what a user might post.
@@ 3 Experimental setup
We replicate the experimental setup of Bagrow et al. (2019) to measure the predictability of users’ online speech using large language models. We define predictability (or rather unpredictability) as a measure of the model’s uncertainty, i.e., its negative log-likelihood, on a specific user’s tweets. We observe how model uncertainty changes given additional sources of context, specifically:
  1. 1.
    user context: past tweets of the user
  2. 2.
    peer context: tweets from the user’s social circle
  3. 3.
    random context: tweets from randomly selected users
We start by describing our data collection process (Section 3.1) and what models we used (Section 3.2). Then, we go into detail about the implementation of our prompting and finetuning experiments (in Section 3.3 and 4.3, respectively). We don’t share the tweets we collected because of X’s Developer Policy. Our code can be found here: https://github.com/socialfoundations/twitter-predictability
@@
Figure 3: Our data collection process can be divided into two stages. In the first stage (Fig. 3(a)), we collected 10M tweets in early 2023 which served as our base for sampling subjects. In the second stage (Fig. 3(b)), we collected users’ timelines.
@@ 3.1 Data Collection
The data collection process consisted of two main stages: an initial sampling period where we recorded real-time Twitter activity for a month in early 2023 and a second stage where we collected the timelines of sampled users. For a high-level overview of the data collection process, see Figure 3. We collected tweets from three groups of users: (i) subjects, approximately 5,000 randomly sampled Twitter users, (ii) peers of subjects, which we take to be the top 15 people that each subject most frequently @-mentions, (iii) and random users for control purposes.
Throughout our data collection, we only collected tweets that were written by the user who posted them (e.g. no retweets) and were classified as English according to Twitter’s own classification algorithm. We additionally preprocessed our tweets (removed urls, special characters, etc.). For more detailed information on the dataset and how it was collected, we refer the reader to Appendix A.1. There we go into detail on what methods we used to achieve a high-quality, representative dataset of English tweets.
3.1.1 Sampling stage
Sampling was done by collecting a pool of tweets whose authors would serve as our base for picking our subjects. We collected them using the Twitter Firehose API; which allowed us to collect a 1% sub-sample of real-time tweet activity. This collection period lasted roughly 30 days and was done in early 2023 (from 20. January to 10. February), during which we collected 10M tweets (with $\sim$5M unique authors). We sampled $N=5102$ subjects from this pool ($0.1\%$ of authors) for our experiments. We filtered out users that scored high (above 0.5 on a scale of 0 to 1) on the Bot-O-Meter bot detection API Sayyadiharikandeh et al. (2020) as well as users that had a high retweet ratio. This ensured the selection of authentic users who had a sufficient amount of self-authored tweets for our dataset.
3.1.2 Timeline collection stage
For each subject, we collected a total of 500 tweets $\mathcal{T}_{u}=\mathcal{T}^{\texttt{eval}}_{u}\cup\mathcal{T}^{\texttt{user}}_{u}$ from their timelines. Half of those tweets were used for estimating predictability $\mathcal{T}^{\texttt{eval}}_{u}$, while the other half $\mathcal{T}^{\texttt{user}}_{u}$ served as context. Besides user context, we also introduce peer and random context: $\mathcal{T}^{\texttt{peer}}_{u}$ and $\mathcal{T}^{\texttt{random}}_{u}$, again with 250 tweets each. Peer context contained tweets from users that the subject most frequently @-mentioned (top-15). We made sure that all context tweets were authored before the oldest tweet in $\mathcal{T}^{\texttt{eval}}_{u}$. In total, we collected approximately 6.25M tweets from users’ timelines.
Similarly to Bagrow et al. (2019), we created a second control group (temporal control). However, because of the similarity of the results, we only report results on the random control group (social control) in the main text. For results including the temporal control, please see Appendix B.5.
3.2 Models
We used four different model families for our experiments, namely GPT-2 Radford et al. (2019), Falcon Almazrouei et al. (2023) as well as Llama-2 Touvron et al. (2023) and Llama-3 Dubey et al. (2024). These represent models that share a similar transformer-based architecture that have recently become popular due to their impressive generative capabilities. However, they differ in number of parameters (ranging between 1.5B-70B), training corpus and tokenizers. We further differentiate between base and finetuned versions of these models:
@@ 3.2 Models
    base: pre-trained LLMs, no instruction-tuning (e.g. GPT-2-XL)
@@
    tweet-tuned: base models finetuned on the 10M tweets collected during the sampling stage (e.g. GPT-2-XL-tt)
For more details on tweet-tuning, please see Appendix A.2. We considered using GPT-3 and GPT-4, however the OpenAI API unfortunately does not allow access to log probabilities for all tokens (which is necessary for our analysis), only to the top-5. We used Huggingface’s transformers library Wolf et al. (2020) to load the models and run our experiments.
@@ 3.3 Prompting Experiments
Our first approach to measuring model uncertainty is through prompting; that is, experiments where we feed a tweet to a model and observe the associated probabilities of outputting that exact tweet. We use negative log-likelihood (NLL) as a measure of model uncertainty, and introduce bits per character (or BPC) to enable comparisons across models. The NLL of a tweet $T=(t_{1},t_{2},...t_{m})$ is commonly defined as follows: $L(T)=-\sum_{t_{i}\in T}\ln{p_{\theta}(t_{i}|t_{<i})},$ where we use a language model with parameters $\theta$ to predict token $t_{i}$ based on the preceding tokens $t_{<i}$. Let $\bar{L}_{u}$ be the average uncertainty associated with predicting tweets of user $u$:
@@
where $n$ is the total number of tokens. A model-agnostic version of this uncertainty is bits per character (BPC), or otherwise known as bits per byte:
@@
where $\bar{C}_{u}$ is the average number of characters per token for user $u$. $\overline{bpc}_{u}$ tells us the average number of bits required to predict the next character of user $u$’s tweets. In the main text we will commonly report the average BPC over all users, which is $\overline{bpc}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\overline{bpc}_{u}$. Reported results are calculated on $\mathcal{T}^{\texttt{eval}}_{u}$.
We additionally introduce notation to distinguish what context was used to calculate: $\overline{bpc}^{c}$, where subscript $c\in\{\texttt{user},\texttt{peer},\texttt{random}\}$. Here, the conditional probability of token $t_{i}$ is based on preceding tokens $t_{<i}$ as well as tokens from the appropriate context: $p_{\theta}(t_{i}|\mathcal{T}_{u}^{c},t_{<i})$. The added context lends a convenient cross-entropy like interpretation of the shared information between the context and evaluation tweets we are trying to predict.
We also quantify the average effect size of different contexts on predictability relative to each other. We do this by calculating the standardized mean difference (SMD) of the negative log-likelihoods calculated using different sources of context, i.e. the SMD of $\Delta^{c1}_{c2}(u)=\bar{L}_{u}^{c1}-\bar{L}_{u}^{c2}$. Aggregating over all users, this gives us a unit-free quantity to estimate the difference in predictability one context offers over another. Following established practice Cohen (1988), we characterize effect sizes up to $0.2\sigma$ to be small, up to $0.8\sigma$ to be medium and and anything above that to be a large effect size. Larger models (>1.5B parameters) were loaded using 8-bit precision, which has little to no impact on performance Dettmers et al. (2022). For more details on why we use BPC, sensitivity to prompting strategy, context size, etc. please refer to Appendix A.3.
@@ 3.4 Finetuning Experiments
In this section we present a second method for estimating predictability, which has three advantages over our previous prompting approach. First, we can fit the entire context into the model compared to prompting where the maximum size of the input sequence is a limitation. Second, it bypasses a common ailment of LLMs: their sensitivity to prompting strategy. Finally, it avoids any limitations in-context learning might have.
Instead of including said context in our prompt, we finetune our model on the different types of context we previously introduced and quantify model uncertainty using its cross-entropy loss in the final round of finetuning. We also include experiments on finetuning on a mixture of contexts (eg. peer+random containing tweets from both $\mathcal{T}^{\texttt{peer}}_{u}$ and $\mathcal{T}^{\texttt{random}}_{u}$). We combine them by sampling an equal amount of tweets uniformly from each context, while keeping the total number of tweets the same (250 tweets) to enable a fair comparison. We finetuned for 5 epochs with constant learning rate $1\mathrm{e}{-5}$ and batch size 1. We tracked the cross-entropy loss on $\mathcal{T}^{\texttt{eval}}_{u}$ periodically. Reported results are calculated on $\mathcal{T}^{\texttt{eval}}_{u}$.